hydrogen may be liquefied at -235 degree celsius under a pressure of 20 atm. the temperature on the degree Fahrenheit scale is...
1
Expert's answer
2021-02-07T19:06:56-0500

We can convert celsius to Fahrenheit using the following formula:


"T_{^{\circ}F}=T_{^{\circ}C}\cdot \dfrac{9}{5}+32,""T_{^{\circ}F}=-235^{\circ}C\cdot \dfrac{9}{5}+32=-391^{\circ}F."